OpenStudy (anonymous):

The cost for 6 dinners plus a 20% tip is $36.00. Write an equation representing the situation if p is the price of the dinner in dollars. You may assume there is no tax.

OpenStudy (anonymous):

@one2345

OpenStudy (anonymous):

p = the price of the dinner in dollars 0.2(36) represent the tip 6p+0.20(36)=$36 please make sure you understand how to get the answer so that you know how to do similar problems!

OpenStudy (anonymous):

its not right :(

OpenStudy (anonymous):

can you write out the equation? @one2345

OpenStudy (anonymous):

what do you mean? 0.2*36 = $7.20 tip 6p+7.20 = 36

OpenStudy (anonymous):

can you see if you know this instead? its saying its wrong. Consider triangle △ABC in the plane where A=(−6,5) B=(−34,21) C=(−4,12) Find the lengths of the sides of the triangle: AB= BC= AC= Is △ABC a right triangle? ( yes or no )

OpenStudy (anonymous):